# Yue Liao  Pengfei Zhou  Siyuan Huang  Donglin Yang  Shengcong Chen Yuxin Jiang   Yue Hu  Jingbin Cai   Si Liu   Jianlan Luo   Liliang Chen Shuicheng Yan   Maoqing Yao   Guanghui Ren†⋄

### v2

@@ Introduction
To navigate the trade-off between efficiency and capacity in robotic video modeling, we utilize a compact video generation model as the core architecture. Our GE-Base world model $\mathcal{W}$ is designed with flexibility in mind, enabling seamless integration with a variety of diffusion transformer (DiT)-based video generation models, e.g., LTX-Video 2B (HaCohen et al., 2024) and COSMOS 2B (Agarwal et al., 2025). Considering the egocentric nature of perception in dual-arm robotic systems, we extend $\mathcal{W}$ to a multi-view, language-and-image-conditioned generation framework that leverages temporally synchronized inputs from three on-board cameras: a head-mounted view ($v^{h}$), and two wrist-mounted views ($v^{l}$, $v^{r}$). Each frame in $x_{0}$, $\hat{\mathbf{x}}_{0:t-1}$, and $x_{t}$ comprises this tri-view observation structure.
@@
Stage I: Multi-Resolution Temporal Adaptation (GE-Base-MR). The first stage aims to bridge the gap between generic video representation learning and robotic-specific motion dynamics. We pretrain the model on 57-frame video sequences randomly sampled at frame rates between 3 Hz and 30 Hz. Each training sample includes four sparse memory frames, randomly drawn from prior video history to augment temporal diversity. These clips are encoded into an 8-frame latent space via a pretrained VAE, where noise is added and the model is optimized through a denoising objective.
This training setup exposes the model—termed GE-Base-MR—to a broad spectrum of motion speeds and temporal patterns, encouraging it to learn spatiotemporal representations that are invariant to sampling rates. Conditioning on both visual conditions and language instructions, the model is trained to map high-level task intent to low-level visual dynamics while maintaining robustness to partial observations. This design is critical for real-world deployment, where sensor latency, frame drops, and asynchronous data are common. GE-Base-MR is trained end-to-end on the AgiBot-World-Beta dataset using 32 NVIDIA A100 GPUs over a period of approximately 7 days.
@@
Bridging high-level world modeling and low-level control is essential for deploying vision-language foundation models in embodied robotics. We present GE-Act, a plug-and-play world action module that augments the GE-Base foundation model with a lightweight, 160160M-parameter autoregressive action decoder. GE-Act translates multimodal latent representations—conditioned on multi-view visual observations and language instructions—into temporally structured action policies, enabling instruction-following behavior without explicit video generation. This architecture supports tight coupling between perception and control, offering a scalable and efficient solution for real-time robotic manipulation in diverse environments.
@@
We realize this capability by extending the GE-Base foundation model into an action-conditioned world simulator, GE-Sim. In this formulation, action trajectories serve as the primary control signal for driving video synthesis over time. To promote visual consistency across generated frames, we incorporate a reference image, encoded by a frozen CLIP image encoder, as a lightweight style anchor. This image is injected via cross-attention into each DiT block, complementing the retained visual observations that provide spatial grounding throughout the sequence.
@@
To assess the precision of action-conditioned video generation, we visualize simulation outputs from GE-Sim based on ground-truth control sequences. As shown in Figure 15, each example displays the current observation frame overlaid with the projected target position of the next action, alongside the corresponding predicted frame synthesized by the simulator. Across tasks and viewpoints, the generated end-effector motion consistently aligns with the spatial intent of the action input, demonstrating that GE-Sim can faithfully and precisely translate low-level control commands into coherent visual predictions. This result further validates the reliability of GE-Sim as a high-fidelity, action-aligned video world simulator for robotic manipulation.
@@
Temporal alignment (TA) is then evaluated using Normalized Dynamic Time Warping (NDTW) , which captures consistency in both sequence and timing between the generated and ground truth trajectories. To produce a positively correlated metric, we report the inverse of the NDTW distance:
@@
As shown in Figure 16, GE-Base consistently outperforms the baselines across multiple evaluation dimensions, with notable strengths in temporal alignment and dynamic consistency—two core metrics for generating action-plausible and temporally stable robotic behaviors. While performance in motion semantics is comparable to generic video generation models, GE-Base demonstrates much stronger control-aware generation fidelity, offering more precise and reliable task execution. This advantage is attributed to GE-Base’s pretraining on large-scale robotic manipulation data, which better equips it to capture task-relevant spatial-temporal dynamics.
@@
In addition to instruction-conditioned evaluation, we assess the fidelity and reliability of our video-based simulator in an action-conditioned setting. Given ground-truth action trajectories, the simulator generates visual predictions conditioned solely on these control sequences. Under the proposed EWMBench framework (Table 2), GE-Sim consistently achieves high alignment in spatial configuration, trajectory execution, and semantic coherence. The low visual diversity observed under fixed action inputs reflects a desirable property, indicating precise action-to-video correspondence and strong alignment with control dynamics.
@@
BLEU CLIP DYN Diversity
0.3255 90.79 0.7836 0.0111
SA Logic TA Scene
0.9413 0.9667 0.9750 0.9016
Table 2: GE-Sim Evaluated with EWMBench. We assess the quality of action-conditioned video generation using our EWMBench framework, measuring spatial, temporal, and semantic alignment with ground-truth control trajectories.
@@
To validate the reliability and task relevance of our proposed EWMBench, we conduct a comparative analysis against human preference ratings and the general video benchmark VBench. We collect human annotations on videos generated by four representative models—GE-Base, Kling-1.6 (Kuaishou, 2025), Hailuo I2V-01-live (MiniMax, 2024), and OpenSora-2.0 (Zheng et al., 2024)—using a ranking protocol where annotators assign ordinal scores based on perceived overall quality. Rankings are aggregated across annotators and samples, with multiple review rounds to ensure annotation consistency. As shown in Figure 17, empirical results demonstrate that EWMBench rankings exhibit strong concordance with human judgments, effectively capturing dimensions of temporal alignment, semantic fidelity, and visual coherence. In contrast, VBench exhibits misalignment, particularly in scenarios demanding embodied consistency and goal-conditioned reasoning. These results confirm that EWMBench provides a more faithful and task-grounded assessment of video-based world models in robotic manipulation.

### v3

@@ Introduction
To balance efficiency and capacity in robotic video modeling, we adopt a compact video generation model as the core architecture. Our GE-Base world model $\mathcal{W}$ is designed with flexibility in mind, allowing seamless integration with various diffusion transformer (DiT)-based video generation models. Specifically, we select LTX-Video 2B (HaCohen et al., 2024) and COSMOS2 2B (Agarwal et al., 2025) as our base models. LTX-Video provides a faster and more lightweight architecture, supporting efficient downstream action policy prediction, whereas COSMOS2 offers higher-quality video synthesis, making it well-suited for high-fidelity simulation. Considering the egocentric nature of perception in dual-arm robotic systems, we extend $\mathcal{W}$ into a multi-view, language-and-image-conditioned generation framework that leverages temporally synchronized inputs from three onboard cameras: a head-mounted view ($v^{h}$) and two wrist-mounted views ($v^{l}$, $v^{r}$). Each frame in $x_{0}$, $\hat{\mathbf{x}}_{0:t-1}$, and $x_{t}$ follows this tri-view observation structure.
@@
Stage I: Multi-Resolution Temporal Adaptation (GE-Base-MR). The first stage aims to bridge the gap between generic video representation learning and robotic-specific motion dynamics. We pretrain the model on 57-frame video sequences randomly sampled at frame rates between 3 Hz and 30 Hz. Each training sample includes four sparse memory frames, randomly drawn from prior video history to enhance temporal diversity. These clips are encoded into an 8-frame latent space via a pretrained VAE, where noise is added and the model is optimized through a denoising objective.
This training setup exposes the model, referred to as GE-Base-MR, to a wide spectrum of motion speeds and temporal patterns, encouraging it to learn spatiotemporal representations invariant to sampling rates. Conditioning on both visual observations and language instructions, the model learns to map high-level task intent to low-level visual dynamics while maintaining robustness to partial observations. This design is essential for real-world deployment, where sensor latency, frame drops, and asynchronous data are common. After this stage, GE-Base-MR is capable of generating high-quality robotic manipulation videos that accurately capture motion dynamics and visual consistency. The model is trained end-to-end on the AgiBot-World-Beta dataset using 32 NVIDIA A100 GPUs for approximately seven days.
@@
Bridging high-level world modeling and low-level control is essential for deploying vision-language foundation models in embodied robotics. We present GE-Act, a plug-and-play world action module that augments the fast LTX-Video–based GE-Base foundation model with a lightweight 160M-parameter autoregressive action decoder. GE-Act translates multimodal latent representations—conditioned on multi-view visual observations and language instructions—into temporally structured action policies, enabling instruction-following behaviors without explicit video generation. This architecture tightly couples perception and control, providing a scalable and efficient solution for real-time robotic manipulation across diverse environments.
@@
We realize this capability by extending the GE-Base foundation model into an action-conditioned simulator, GE-Sim. In this framework, action trajectories serve as the primary control signals driving video synthesis over time. To implement GE-Sim, we adopt two GE-Base architectures: the fast LTX-Video–based variant used in GE-Act, and the COSMOS2 2B–based variant for high-fidelity simulation and realistic video generation. To maintain visual consistency across generated frames, we incorporate a reference image encoded by a frozen CLIP image encoder as a lightweight style anchor. This reference is injected via cross-attention into each DiT block, complementing the spatial grounding provided by visual observations.
@@
Figure 16: Comparison of Multi-View Action-Conditioned Generation Results between Two GE-Sim Variants. Visualization of GE-Sim outputs based on different base models: COSMOS2 and LTX-Video, under identical action conditions.
@@
To evaluate the precision of action-conditioned video generation, we visualize simulation outputs from GE-Sim based on ground-truth control sequences. As shown in Figure 15, each example presents the current observation frame overlaid with the projected target position of the next action, along with the corresponding predicted frame synthesized by the simulator. Across tasks and viewpoints, the generated end-effector motion consistently aligns with the spatial intent of the action input, demonstrating that GE-Sim can accurately translate low-level control commands into coherent visual predictions. Furthermore, we compare GE-Sim built on two base architectures under identical action-conditioning in Figure 16. The COSMOS2-based variant exhibits higher visual fidelity and stronger temporal consistency than the LTX-Video–based model, confirming its superior capability in generating high-quality, action-aligned robotic simulations.
@@
Temporal alignment (TA) is then evaluated using Normalized Dynamic Time Warping (NDTW) (Ilharco et al., 2019), which captures consistency in both sequence and timing between the generated and ground truth trajectories. To produce a positively correlated metric, we report the inverse of the NDTW distance:
@@
As shown in Figure 17, GE-Base consistently outperforms the baselines across multiple evaluation dimensions, with notable strengths in temporal alignment and dynamic consistency, two core metrics for generating action-plausible and temporally stable robotic behaviors. While performance in motion semantics is comparable to generic video generation models, GE-Base demonstrates much stronger control-aware generation fidelity, offering more precise and reliable task execution. This advantage is attributed to GE-Base’s pretraining on large-scale robotic manipulation data, which better equips it to capture task-relevant spatial-temporal dynamics.
@@
In addition to instruction-conditioned evaluation, we further assess the fidelity and reliability of our video-based simulator in an action-conditioned setting across two base models. Given ground-truth action trajectories, the simulator generates visual predictions conditioned solely on these control sequences. Under the EWMBench framework (Table 2), GE-Sim consistently demonstrates high spatial accuracy, precise trajectory execution, and strong semantic coherence. The limited visual diversity under fixed action inputs indicates accurate action-to-video correspondence and robust alignment with control dynamics. While GE-Sim built on COSMOS2 and LTX-Video achieves comparable overall performance, the COSMOS2-based variant excels in dynamic consistency, reflecting superior embodied video generation fidelity and temporal coherence.
@@
Model BLEU CLIP DYN Div. PSNR
LTX 0.33 90.8 0.78 0.011 19.9
COSMOS 0.31 90.2 0.85 0.010 20.7
Model SA Log. TA Scn.
LTX 0.94 0.97 0.98 0.90
COSMOS 0.87 0.97 0.97 0.91
Table 2: GE-Sim Evaluation on EWMBench. Comparison of LTX-Video and COSMOS2 as base models for action-conditioned video generation. Metrics evaluate spatial, temporal, and semantic alignment with ground-truth control trajectories.
@@
To validate the reliability and task relevance of our proposed EWMBench, we conduct a comparative analysis against human preference ratings and the general video benchmark VBench. We collect human annotations on videos generated by four representative models, GE-Base, Kling-1.6 (Kuaishou, 2025), Hailuo I2V-01-live (MiniMax, 2024), and OpenSora-2.0 (Zheng et al., 2024), using a ranking protocol where annotators assign ordinal scores based on perceived overall quality. Rankings are aggregated across annotators and samples, with multiple review rounds to ensure annotation consistency. As shown in Figure 18, empirical results demonstrate that EWMBench rankings exhibit strong concordance with human judgments, effectively capturing dimensions of temporal alignment, semantic fidelity, and visual coherence. In contrast, VBench exhibits misalignment, particularly in scenarios demanding embodied consistency and goal-conditioned reasoning. These results confirm that EWMBench provides a more faithful and task-grounded assessment of video-based world models in robotic manipulation.